An 30 kg object from rest was pulled horizontally by a pulling force of 100N applied along the direction of displacement line. The coefficient of kinetic friction between the surface and the object is 0.20 and the object was displaced at a distance of 40 meters. How work is done on friction?



1
Expert's answer
2021-10-20T10:42:41-0400

By the definition of the work done we have:


"W_{fr}=F_{fr}dcos\theta=\mu_kNdcos\theta=\mu_kmgdcos\theta,""W_{fr}=0.2\times30\ kg\times9.8\ \dfrac{m}{s^2}\times40\ m\times cos180^{\circ}=-2352\ J."

The sign minus means that the force of friction acts in the opposite direction to the motion of the object and, therefore, does negative work.